Polynomial and rational functions can be used to model a wide variety of phenomena of science, technology, and everyday life.

Choose one of these sectors and give an example of a polynomial or rational function modeling a situation in that sector. [Hint: see the examples and exercises in the book.]


1
Expert's answer
2021-07-05T06:51:41-0400

People use polynomials for modeling various buildings and objects , used in industries, construction, marketing , finance , stocks and many other areas. Now it is common that civil engineer use polynomials to design roads, buildings and other structures.

an example of polynomial and rational function

function f(x)= (3x-2)(x+2)2

y intercept =

f(0)= {3(0)-2}(0+2)2

= (-2)(4)

= -8

y intercept of the graph of y=f(x) is(0,-8)


x intercept

0= (3x-2)(x+2)2

3x-2=0 , x=2/3

(x+2)2=0 , x=-2

x intercepts of the graph of y=f(x) are (2/3,0) (-2,0).

Now find the end behavior

f(x)= (3x-2)(x+2)2

f(x)= (3x-2)(x2+4x+4)

f(x) = 3x3+12x2+12x-2x2-8x-8

f(x) =3x3+10x2+4x-8

 the end behavior of function f will be the same as the end behavior of 3x3.